Angelito is driving at 6.0 m/s when a traffic enforcer signals that he should speed up. He then steps on the gas pedal and accelerated uniformly to a speed of 15 m/s in 4.5 s. Find the total distance Angelito traveled. Assume Angelito's reaction time is 0,15 s.
1
Expert's answer
2020-11-30T14:59:49-0500

"v=v_0+at\to a=\frac{v-v_0}{t}=\frac{15-6}{4.5}=2(m/s^2)"


"S=\frac{v^2-v_0^2}{2a}=\frac{15^2-6^2}{2\cdot 2}=47.25(m)"


"S_0=0.15\cdot6=0.9(m)"


"S_{total }=S+S_0=0.9+47.25=48.15(m)"